OpenStudy (nightmarenight):

Please help? Calculate the discriminant. x^2 – 2x + 4 = 0

OpenStudy (agent0smith):

\[\Large ax^2 + bx + c = 0\] has a discriminant \[\Large b^2 - 4ac\]

OpenStudy (anonymous):

The discriminant is\[b^2-4ac\] For reference, the equation for a parabola is \[y=ax^2+bx+c\]

OpenStudy (nightmarenight):

okay, I dunno what to do next @agent0smith

OpenStudy (agent0smith):

Well, identify what a, b and c are... and plug them into the discriminant.

OpenStudy (nightmarenight):

What? if I take what changed it doesnt fit

OpenStudy (anonymous):

If the discriminant is greater than 0, there are 2 real roots. If it is equal to 0, it has 1 real and 1 complex root. If it is less than 0, it has 2 complex roots.

OpenStudy (nightmarenight):

this isn't making any sense to me, I'm sorry I'm not catching on @halorazer

OpenStudy (anonymous):

Hmm. I will steal an example from a math website. in y = 3x² + 9x + 5 You would put the variables in. a=3 b=9 c=5 Then in the discriminant, which is\[b^2-4ac\] \[(9)^2-4(3)(5)\] Then we simplify it. \[81-60=21\] Because 21 is greater than zero, there are two real roots. If it was equal to exactly zero, there would be 1 real root and 1 complex one. If it was negative, there would be 2 complex roots.

OpenStudy (agent0smith):

"What? if I take what changed it doesnt fit" don't know what this means, but for x^2 – 2x + 4 a = 1 b = -2 c = 4 Plug them into the discriminant.

OpenStudy (nightmarenight):

x^2 – 2x + 4 = 0 changes to variables. Right? @agent0smith

OpenStudy (agent0smith):

Just compare it to ax^2 + bx + c to identify each.

OpenStudy (nightmarenight):

Where did you get a=1?

OpenStudy (nightmarenight):

there's no 1.

OpenStudy (agent0smith):

What's the coefficient of x^2? It's not zero...

OpenStudy (nightmarenight):

unless you changed x

OpenStudy (nightmarenight):

Ah.

OpenStudy (agent0smith):

The coefficient of x, or y, or x^2... is 1.

OpenStudy (nightmarenight):

okay, I got a-bx+c????????

OpenStudy (nightmarenight):

@agent0smith ^^

OpenStudy (agent0smith):

I have no idea what you are doing... We already know what a, b and c are from above: a = 1 b = -2 c = 4 Now plug those numbers into \[\Large b^2 - 4ac\]